# Asymptotic Behavior of Rupture Solutions for the Elliptic MEMS Equation with Hénon-Type Term

### v1

@@ 4 Existence of Solutions
$\mathcal{N}(\eta_{\ast}+\phi)=\mathcal{L}(\phi)-\Big[(\eta_{\ast}+\phi+\Lambda)^{-p}-\Lambda^{-p}+p\Lambda^{-(p+1)}(\eta_{\ast}+\phi)\Big]=\mathcal{L}(\phi)-\sum_{i=2}^{\infty}b_{i}(\eta_{\ast}+\phi)^{i}.$ (71)
@@
We now write the argument for case (31) in the proof of Theorem 2.2. The associated index sets $\mathcal{I}_{\rho}$ and $\mathcal{I}_{\widetilde{\rho}}$ are those given in (37) and (38) (ignoring multiple numbers). In this case,
@@
$\mathcal{N}\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l}\big)=O\big(e^{\widetilde{\rho}_{l+1}t}\big).$ (74)
@@
$\mathcal{N}(\eta_{\ast}+\varphi)=\sum_{n_{1}+\cdots+n_{r_{1}}\geq 2}a_{n_{1}\ldots n_{r_{1}}}e^{(n_{1}\rho_{1}+\cdots+n_{r_{1}}\rho_{r_{1}})t}Q_{1}^{n_{1}}\cdots Q_{j}^{n_{j}}Q_{0}^{n_{j+1}}Q_{j+1}^{n_{j+2}}\cdots Q_{r_{1}-1}^{n_{r_{1}}},$
@@
$\mathcal{N}(\eta_{\ast}+\varphi)=\sum_{i=1}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{K+1}t}\big),$ (75)
@@
$\mathcal{N}\left(\eta_{*}+\varphi\right)=O\left(e^{\tilde{\rho}_{1}t}\right),$
@@
$\mathcal{N}\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l}\big)=\mathcal{L}(\widetilde{\varphi}_{1})+\cdots+\mathcal{L}(\widetilde{\varphi}_{l})+\sum_{i=1}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big),$
@@
$\mathcal{N}\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l-1}\big)=\sum_{i=l}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big),$
@@
$\mathcal{N}\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l}\big)=\mathcal{L}(\widetilde{\varphi}_{l})+\sum_{i=l}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big).$
@@
$\mathcal{N}(\eta_{\ast}+\varphi+\widetilde{\varphi})=O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big)=O(e^{\mu t}).$
@@
where $c_{ijm}$ are constants. This completes the proof of the proposition.
@@
we obtain the desired solution $u$, and the proof of Theorem 1.1 and Thereom 1.2 is thereby completed.
@@ Acknowledgments
Conflict of interest

### v3

@@ 4 Existence of Solutions
$H(\eta_{\ast}+\phi)=\mathcal{L}(\phi)-\Big[(\eta_{\ast}+\phi+\Lambda)^{-p}-\Lambda^{-p}+p\Lambda^{-(p+1)}(\eta_{\ast}+\phi)\Big]=\mathcal{L}(\phi)-\sum_{i=2}^{\infty}b_{i}(\eta_{\ast}+\phi)^{i}.$ (68)
@@
We now write the argument for case (31) with $k=1$ in the proof of Theorem 2.2. The associated index sets $\mathcal{I}_{\rho}$ and $\mathcal{I}_{\widetilde{\rho}}$ are those given in (34) and (35) (ignoring multiple numbers). In this case,
@@
$H\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l}\big)=O\big(e^{\widetilde{\rho}_{l+1}t}\big).$ (71)
@@
$H(\eta_{\ast}+\varphi)=\sum_{n_{1}+\cdots+n_{r_{1}}\geq 2}a_{n_{1}\ldots n_{r_{1}}}e^{(n_{1}\rho_{1}+\cdots+n_{r_{1}}\rho_{r_{1}})t}Q_{1}^{n_{1}}\cdots Q_{j}^{n_{j}}Q_{0}^{n_{j+1}}Q_{j+1}^{n_{j+2}}\cdots Q_{r_{1}-1}^{n_{r_{1}}},$
@@
$H(\eta_{\ast}+\varphi)=\sum_{i=1}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{K+1}t}\big),$ (72)
@@
$H\left(\eta_{*}+\varphi\right)=O\left(e^{\tilde{\rho}_{1}t}\right),$
@@
$H\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l}\big)=\mathcal{L}(\widetilde{\varphi}_{1})+\cdots+\mathcal{L}(\widetilde{\varphi}_{l})+\sum_{i=1}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big),$
@@
$H\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l-1}\big)=\sum_{i=l}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big),$
@@
$H\big(\eta_{\ast}+\varphi+\widetilde{\varphi}_{0}+\cdots+\widetilde{\varphi}_{l}\big)=\mathcal{L}(\widetilde{\varphi}_{l})+\sum_{i=l}^{\widetilde{K}}\left\{\sum_{m=0}^{\widetilde{M}_{i}}a_{im}Q_{m}(\theta)\right\}e^{\widetilde{\rho}_{i}t}+O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big).$
@@
$H(\eta_{\ast}+\varphi+\widetilde{\varphi})=O\big(e^{\widetilde{\rho}_{\widetilde{K}+1}t}\big)=O(e^{\mu t}).$
@@
where $c_{ijm}$ are constants. This completes the proof of the proposition. ∎
@@
we obtain the desired solution $u$, and the proof of Theorem 1.1 and Theorem 1.2 is thereby completed.
@@ Acknowledgments
Y. Zhang is sponsored by NSFC [No. 12271505] and STCSM [No. 22DZ2229014].